### v1

# %First␣names␣are␣abbreviated␣in␣the␣running␣head.%If␣there␣are␣more␣than␣two␣authors,␣’et␣al.’␣is␣used.http://www.docugami.com

Abstract
Mathematical reasoning capabilities are increasing with tool-augmented language agents, but methods often rely either on closed-source or large models, external data, or extensive prompt engineering. This work introduces MATATA, a novel cost-effective method to train LLM agents for tabular data problems through reasoning, planning, and tool use. With a progressive self-improvement paradigm and an iterative weak supervision, it empowers 3.8B/8B Small Language Models (SLMs), particularly suited for local hosting and sensitive business contexts where data privacy is crucial. By employing a flexible and reusable tools across different datasets, it achieves robust performance with effective scalability across shared tasks. Experiments show that MATATA reaches state-of-the-art performances on FinQA and TAT-QA among reasoning frameworks based on open-source models. Moreover, MATATA models compete with GPT-4 based frameworks on TabMWP, while being SLMs.
Keywords:
Mathematical Reasoning Small Language Model Weak-Supervision.
1 Introduction
Recent developments have popularized Large Language Models (LLMs) [1, 24, 3] as potent tools for tackling complex mathematical reasoning problems [18]. Different strategies encompass an extensive range of techniques, including specialized models like [21, 31], Chain-of-Thought or Program-of-Thought [25, 4], or, tool-based compositional frameworks like [8, 14, 16, 27]. Progress in the study of mathematical reasoning with structured and unstructured data has been tangible thanks to various datasets and benchmarks, such as [5, 30, 17]. Tool-augmented LLMs are particularly prevalent in those benchmarks as they combine program assistance, data manipulation, natural language reasoning, and many other tools. These frameworks leverage the increasing planning capabilities of LLMs [26, 2] to orchestrate the tools, enabling multi-step problem-solving.
Figure 1: Example from FinQA with questions over tabular and textual data.
Some studies focus on prompt engineering methods to get the best from closed-source models like GPT-4 [22, 16, 25, 4]. Others involve fine-tuning multiple specialized Small Language Models (SLMs) in a teacher-student fashion with data annotated automatically by larger models [14, 8], or extensively by human experts [31, 30, 5]. Other methods focus on self-improvement, with either training verifiers [10], generators for data augmentation [28], or multi-sampling techniques [29]. The diversity in the types of approaches underscores the potential of language models in enhancing reasoning capabilities.
However, each approach faces dependencies or limitations, summarized in Table 1. Prompt engineering is labor-intensive, usually requires hand-crafting efforts, and may not fully capture data complexity and diversity. Closed-source models raise privacy concerns, especially with business documents, and present scalability issues due to their cost and latency, whether for querying models directly or generating high-quality reasoning data. Moreover, using teacher models to generate reasoning steps is often data-inefficient [10], discarding incorrect solutions, and human expert annotation [31] is a costly alternative. To overcome this issue, various self-improvement methods have been developed. However, most of them require multiple samplings and/or additional models [13, 10], leading to an increase in computational and time costs.
Inspired by those challenges, this work introduces a novel approach for mathematical reasoning over tabular data. The proposed method aims to reduce both prompt engineering efforts, reliance on larger or external models, privacy concerns, and computational inefficiencies while maintaining comparable performances to existing methods.
Table 1: Overview of LLM reasoning frameworks over tabular data. PE: Prompt Engineering; FT: Fine-Tuning.
Method Tool-use GPT-4 reliance Human-experts Datasets used
[…]
The main contributions of this work are: 1) a tool-augmented framework that composes shared tools from diverse datasets; 2) a weak-supervised fine-tuning for the tools from self-generated correct reasoning trajectories; 3) a weak-supervised preference optimization to align the tools using all updated reasoning trajectories.
2 Related Work
2.1 Datasets for Mathematical Reasoning
Evaluating LLMs’ mathematical reasoning capabilities is usually focused on datasets like MATH [9] and GSM8K [6], which explore mathematical problem-solving. While these benchmarks span a broad domain of mathematical subjects, the data format is limited to textual word problems. In contrast, other benchmarks like FinQA [5] and TAT-QA [30] represent more complex and real-world reasoning scenarios. These datasets include tabular and textual data, requiring reasoning over financial reports with intricated contexts. Moreover, TabMWP [17] expands to mathematical word problems over different subjects. Those three datasets need multi-step reasoning, information extraction, data manipulation, tabular and contextual understanding and numerical calculations.
2.2 Tool-Augmented Frameworks
Multi-step reasoning has emerged as an approach for mathematical reasoning, particularly over tabular data. Diverse prompting techniques [25, 4, 15, 22] have been explored to encourage stepwise and critical thinking, traditionally relying on closed-source LLMs. Other frameworks adopt LLMs as tools with a dedicated planning: both TORA [8] and Husky [14] fine-tune open-source LLMs with training data generated by GPT-4, while Chameleon [16] relies on prompt engineering, directly backed up by ChatGPT/GPT-4. TAT-LLM [31] relies on a fixed stepwise pipeline for reasoning decomposition, with a fine-tuned model trained on data heavily annotated by human experts. Those fine-tuning approaches rely on closed-source models or human-annotations to generate high-quality synthetic data to train on.
2.3 Alignment Methods
To align LLMs on downstream tasks, methods have progressed from Reinforcement Learning from Human Feedback (RLHF) [19] to Direct Preference Optimization (DPO) [20], and more recently Kahneman-Tversky Optimization (KTO) [7]. DPO offers a cost-effective alternative to RLHF by eliminating the need for a separately trained reward model. Instead DPO relies on a preference dataset with positive and negative completions for given inputs. In contrast to traditional methods that optimize the log-likelihood of preferences, KTO focuses on maximizing generation utility. KTO is also less data-intensive than DPO as it leverages binary preference data, i.e., completions paired with boolean labels indicating the correctness of the completion. Therefore, it is particularly well-suited for downstream tasks following Supervised Fine-Tuning (SFT).
3 MATATA: MAthematical Tool-Assisted reasoning for Tabular Applications
3.1 Motivations
MATATA is a novel cost-effective method for training LLM agents designed to tackle tabular data problems through reasoning, planning, and tool use. This work explores the following questions: Can high performance be achieved while reducing reliance on external models, data, and extensive prompt engineering? Can a tool-augmented SLM framework for mathematical reasoning self-improve through weak supervision?
The underlying assumption is that, even with a lacking baseline generating training data, weak supervision provides sufficient control over each tool to enhance its performance. To answer this, our framework takes inspiration from [8, 14, 16] and employs a planner with specialized tools to generate multi-step reasoning trajectories. By decomposing complex problems into subtasks, each of them is then solved locally with fine-tuned SLMs. This approach is particularly relevant when handling tabular contexts from business documents, where privacy concerns preclude using closed-source LLMs. Unlike other existing fine-tuned tool-augmented LLM approaches like [14, 8], MATATA uses fine-tuned specialized models without relying on GPT-4 to generate training multi-step reasoning trajectories. Prompt engineering efforts are also limited to minimize human involvement and demonstrate the system’s ability to improve and evolve with minimal initial guidance, contrary to [16, 22, 4].
3.2 Proposal
3.2.1 MATATA Compositional Reasoning
The compositional reasoning process is described in Algorithm 1. The planning draws inspiration from Chameleon [16]. Given a context $c$ with tabular and textual data, a question $q$, and a pre-defined set of tools $T$ with tool-specific prompts $p_{T}$, the planner decomposes the task into a sequence of sub-tasks. Each sub-task is then executed sequentially with the appropriate specialized tool (program generation, row or column selection, text extraction, etc.). Each tool interacts with $c$ and the previous outputs to generate intermediate results and update the trajectory. Finally, a parsing tool extracts the answer from the last intermediate result of the trajectory.
Algorithm 1 Compositional reasoning
1:context $c$, question $q$, models $(\mathcal{M}_{t})_{t\in T}$, tool-specific instructions $(p_{t})_{t\in T}$, set of tools $T$
2:Trajectory $\tau$ contains the solution
3:$\tau\leftarrow c\oplus q$triangleright Initialize trajectory
[…]
5:for each tool $T_{i}$ in plan $\mathcal{P}$ do
6:     $o_{i}\sim\mathcal{M}_{t}(\cdot|\tau\oplus p_{T_{i}})$triangleright Tool output
7:     $r_{i}\leftarrow\mathcal{E}(T_{i},o_{i})$triangleright Tool execution $\mathcal{E}$
8:     $\tau\leftarrow\tau\oplus\mathcal{U}(c,r_{i})$triangleright Trajectory update $\mathcal{U}$
9:end for
10:return $\tau$
3.2.2 MATATA Weak-Supervision Stages
In the initial stage, tools are few-shot prompted [3] with tool-specific demonstrations. Those prompts contain examples of how each tool should behave but have not been over-engineered, thus reducing human involvement. This baseline system generates reasoning trajectories used as training data. Consequently, the model improves by learning from its own generated reasoning, as shown in [29]. Once the tools are fine-tuned, the tool-specific prompts are replaced by just the tool’s name, mechanically accelerating the inference process, as models have fewer input tokens to process.
A weak-supervised fine-tuning is then applied to the trajectories, to compensate for the limited guidance put into prompt engineering. Only the correctness of the final answer is used to refine the tools without knowing how they should behave, hence the weak signal. Each trajectory is broken down into input/output pairs for the tools involved.
Our work redefines the teacher-student paradigm as the same model generates the trajectories used for weak supervision fine-tuning. The first training stage is SFT [19], where each tool is fine-tuned on prompt/completion pairs from valid trajectories generated by the baseline system. After the SFT stage, tools are aligned with KTO [7]. Using KTO is motivated by weak supervision, as it only requires a binary signal (correct/incorrect). KTO does not require multiple samplings for a given problem, unlike DPO [20], or additional models like verifiers or generators [10], or bootstrapping methods [29]. At tool level, for each prompt/completion pair, the preference label is the weak signal from the final answer, without requiring step-by-step annotations. Using all invalid trajectories during the alignment phase could help increase domain knowledge in areas uncovered with SFT, countering the tendency of other methods to improve mostly in the framework’s areas of proficiency.
This work employs self-improvement through multi-stage fine-tuning with the same model, inspired by [12]. For each tool $t\in\mathcal{T}$, prompt/completion pairs $(\hat{x}^{t},\hat{y}^{t})$ from self-generated trajectories serve to fine-tune LoRA adapters [11]. For SFT, the dataset $\mathcal{D}^{t}_{\text{SFT}}$ consists of pairs extracted from valid trajectories. The negative log-likelihood objective is defined as:
$\mathcal{L}_{\text{SFT}}(M^{t})\ =\ -\mathbb{E}_{(\hat{x}^{t},\hat{y}^{t})\sim\mathcal{D}^{t}_{\text{SFT}}}\sum_{k=1}^{K}\log M^{t}(\hat{y}^{t}_{k}\mid\hat{y}^{t}_{<k},\hat{x}^{t})$ (1)
After the first training stage, the updated tools generate trajectories for the KTO alignment stage. For each tool $t\in\mathcal{T}$, the dataset $\mathcal{D}^{t}_{\text{KTO}}$ includes pairs for correct and incorrect solutions. The KTO objective is defined as in [7]:
$\mathcal{L}_{\text{KTO}}(\pi_{\theta}^{t},\pi_{\text{ref}}^{t})\ =\ -\mathbb{E}_{(\hat{x}^{t},\hat{y}^{t})\sim\mathcal{D}^{t}_{\text{KTO}}}[\lambda_{\hat{y}^{t}}-v(\hat{x}^{t},\hat{y}^{t})]$ (2)
with $v$ the value function, $\lambda$ the loss-aversion coefficient, $\pi_{\text{ref}}^{t}$ the model before alignment, and $\pi_{\theta}^{t}$ the model being aligned. Once fine-tuned and aligned, the framework can be evaluated.
4 Experiments
4.1 Training Details
Experiments are conducted on FinQA [5], TAT-QA [30], and TabMWP [17]. The selected tools for this experiment are inspired by Chameleon [16]. The following tools are shared across datasets: Program Generator, Solution Generator (Natural Language reasoner), Context Extractor, Row/Column Extractor, and Knowledge Retrieval. The task-specific tools are: a unique Planner for each dataset, and a Span Extractor for TAT-QA only. This modular design enables flexible handling of diverse datasets, allowing to add and share new tools when incorporating additional datasets.
Table 2: Accuracy/Exact-Match scores across evaluation datasets. PE: Prompt Engineering; FT: Fine-Tuning. Best overall scores underlined, best section scores bolded.
Framework Model Method FinQA TAT-QA TabMWP
Acc. EM Acc.
Closed-source models
Chameleon [16] ChatGPT PE - - 93.28
GPT-4 PE - - 98.78
EEDP [22] ChatGPT PE 61.88 79.73 -
GPT-4 PE 76.05 88.67 -
PoT [4] PoT-SC-Codex PE 68.1 70.2 81.8
Open-source models
FinQANet [5] RoBERTa FT 61.24 - -
TagOP [30] RoBERTa FT - 50.1 -
TORA [8] Code-LLama-34B FT - - 70.5
Llama2-70B FT - - 74.0
Husky[14] Llama2-7B FT 20.8 42.2 77.6
Llama3-8B FT 20.9 42.3 76.6
TAT-LLM [31] Llama2-7B FT 65.13 76.4 -
Llama2-13B FT 71.93 77.51 -
Llama2-70B FT 76.81 81.42 -
MATATA (Ours) Phi3-mini 3.8B FT 70.10 74.44 96.66
Ministral-8B FT 77.59 77.81 98.12
Train sets are used to generate reasoning trajectories and train the tools. The best model is selected with respect to scores on validation sets. Results on the test sets are reported in Table  2 using the datasets metrics: Accuracy or Exact-Match. All datasets and splits are used entirely, as in their respective papers.
Two light-weight language models are used as baseline and for fine-tuning: microsoft/Phi-3-mini-4k-instruct 3.8B [1] and mistralai/Ministral-8B-Instruct-2410 [23]. All trainings are conducted on a single NVIDIA A100 GPU, for a combined total of about 128 GPU hours. Throughout the training process, one adapter is trained for each tool $t$ in the framework $M_{t}\rightarrow M^{SFT}_{t}\rightarrow M^{SFT+KTO}_{t}$. Both SFT and KTO adapter fine-tuning use the following hyper-parameters: a learning rate of 1e-5 and a batch size of 32 over 10 epochs, with a LoRA $r$ of 64, $\alpha$ of 32 and dropout of 0.05. KTO uses the default $\beta$ parameters at 0.1, and task-specific weights are applied to address class imbalance between valid and invalid trajectories.
4.2 Experimental Results
Table 3: Evolution of MATATA-3.8B and MATATA-8B Accuracy and EM over training stages; with one vs. all training set(s).
Method Training dataset(s) FinQA TAT-QA TabMWP
Phi3-mini 3.8B baseline - 47.06 43.05 83.87
     + weak-sup SFT One 62.00 68.19 92.89
All 67.74 71.56 96.36
     + weak-sup KTO One 69.75 71.74 94.05
All 70.10 74.44 96.66
Ministral 8B baseline - 57.11 51.29 80.89
     + weak-sup SFT One 70.97 75.59 89.85
All 75.68 76.19 96.64
     + weak-sup KTO One 74.63 76.36 93.38
All 77.59 77.81 98.12
The experimental results in Table 2 show the performances of MATATA-Phi3-mini 3.8B and MATATA-Ministral 8B, respectively referred to as MATATA-3.8B and MATATA-8B. MATATA outperforms Husky, which relies on training data annotated by GPT-4, and PoT-Codex, a GPT-3 variant, on all 3 datasets. Both RoBERTa-based baseline models for FinQA and TAT-QA are also surpassed. On TAT-QA, MATATA-8B ranges between TAT-LLM-13B and TAT-LLM-70B, while TAT-LLM uses human-expert annotations for training. The gap between MATATA-8B and TAT-LLM-70B is of only 3.61%, despite the model size decrease. On FinQA, MATATA-8B is above all methods, and MATATA-3.8B ranges between TAT-LLM-7B and -13B, and between ChatGPT and GPT-4 with EEDP-prompting, while being much smaller than those models. On TabMWP, MATATA-8B exceeds all methods not using GPT-4, ranking 2nd on the TabMWP leaderboard, with only a 0.66% difference behind Chameleon GPT-4. For its range of parameters and size, MATATA competes with models fine-tuned on higher-quality reasoning data, larger, or extensively prompted. This also confirms the effectiveness of our method, intuiting that fine-tuning a larger model with weak-supervision could further improve those scores.
To investigate knowledge transfer across datasets with shared tools, two training strategies are explored: (1) training the tools on separate datasets, and (2) training shared tools on the combination of datasets. Results reported in Table 3 show that training tools with weak-trajectories from all datasets outperforms single dataset training, in both SFT and KTO stages. This improvement can be attributed to a better coverage over trajectories of each tool, with the increase and diversification of training data. It also confirms the reusability and scalability of MATATA: the same framework could be enhanced with other datasets and more shared tools, bringing potential improvements to existing results. Also, the performance increase from baseline models to fine-tuned ones is between 12.79% and 31.39% for the 3.8B model, and between 17.23% and 31.39% for the 8B model. Although the baseline model receives minimal initial guidance through low-effort prompt engineering, applying weak-supervision with SFT and KTO enables tools to naturally adopt their expected behaviors, improving the baseline performances significantly.
5 Conclusion
This work explore a novel method for language agents to tackle tabular data problems using only SLMs, without relying neither on larger and/or close-source models, or external data. This is essential for sensitive business documents, such as those in FinQA or TAT-QA, as submitting those to closed-source LLMs can raise security concerns. With minimal efforts spent on hand-crafting the prompts, human labor is also reduced, adapting the framework to its users rather than the opposite. Competitive results validate that tool-augmented SLMs with weak-supervision can meet performances of existing reasoning methods. This work serves as a step for future developments in creating small, reliable, and secure locally-operated SLM frameworks capable of complex reasoning tasks.
{credits}
5.0.1 Acknowledgements
This work was partially supported by Mathematics of Information Technology and Complex Systems (MITACS) Accelerate, grant IT41737.